A rectangle is 6 meters longer than it is wide. The area fo the rectangle is 315 sqaure meters. Find the length.
Found 2 solutions by checkley75, Kari:
Answer by checkley75(3666)

area=l*w
315=x*(x+6)
315=x^2+6x
x^2+6x-315=0
(x-15)(x+21)=0
x-15=0
x=15 answer for the short side
15+6=21 for the long side
proof
15*21=315
315=315

Answer by Kari(1)

area=l*w
315=x*(x+6)
315=x^2+6x
x^2+6x-315=0
(x-15)(x+21)=0
x-15=0
x=15 answer for the short side
15+6=21 for the long side
proof
15*21=315
315=315